During a workout, a person repeatedly lifts a 13-lb barbell through a distance of 1.2 ft . How many "reps" of this lift are required to burn off 160 C ?
Answer #1


Ez= mgh
Ez= the gravitational energy
m= the mass
g= the gravitational acceleration
h= the height.

You need two times the Ez per lift, because it takes energy as well to hold the bell up while taking it down.

13 lb = 13*0.453592 =5.896kg

1.2ft = 1.2*0.3048 = 0.3657m

Energy per lift= 2Ez = 2(mgh) = 2(5.896kg*9.81m/s^2*0.3657m) J = 42.26

1 cal = about 4.19 J

160 cal * 4.19 = 670.4 J

670.4 / 42.26 = 15.864 lifts